# SLOPE and Designing Robust Studies for Generalization

### v1

@@ I.2.3 Proof of Proposition 2
This section details some extended remarks deferred from the main text. Section J.4 discusses defining SLOPE for other types of conditional exchangeability assumptions, including the no unmeasured confounding assumption and the missing at random assumption. Section J.5 comments on the challenge of extending the notion of SLOPE to other bound-based sensitivity models. Section J.6 states the mathematical connection between SLOPE and the marginal interventional effect (Zhou and Opacic,, 2022) with the incremental propensity score intervention (Kennedy,, 2019).
@@
When the functional $\psi(\cdot)$ is vector valued with dimension $p$, the corresponding SLOPE, as defined in Definition 1, is also a vector of $p$ elements. Each dimension of SLOPE represents the robustness of the corresponding element of the target estimand when conditional exchangeability is “near” violated. For example, when $X$ is $p$-dimensional and the target estimand is a $p$-dimensional vector of ordinary least squares (OLS) coefficients of regressing $O$ on $X$, then each element of SLOPE describes the robustness of the corresponding coefficient; see Section C.5 for a formal presentation. More generally, for SLOPE as vectors, the connection between SLOPE and IF still holds (i.e., Theorem 3. In addition, the two proposed estimators are still applicable, as mentioned in Remarks 5 and 6.
The main difference between a vector of SLOPE and a scalar SLOPE is on the interpretation, and in particular, the informed guidance in designs. Specifically, as mentioned in Section 3.1, when SLOPE becomes a vector and all elements are relevant, then practitioners need to find an appropriate summary of the vector in order to compare SLOPEs across study designs. This could be a visual summary, such as plotting the SLOPE of each dimension, or a quantitative measure such as a norm of the SLOPE.
J.2 SLOPE Beyond Exponential Tilting
Remark 7 (The Direction of Sensitivity).
In order to understand the re-centering term $O-\mu(X)$ in equation (10) of Theorem 3, we consider a broader class of sensitivity models that tilts the density ratio by a given non-negative function $\rho(O,X,\gamma)$ and incorporates the exponential tilting model (3) as a special case with $\rho(O,X,\gamma)=\exp(\gamma\cdot O)$:
$\dfrac{dQ_{O\mid X}^{\gamma}}{dP_{O\mid X}}(O,X)$ $\propto\rho(O,X,\gamma),$
where $\rho(O,X,0)=1$ and $\int\rho(O,X,\gamma)dP_{O\mid X}\in(0,\infty)$. Then the result in Theorem 3 becomes:
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\textnormal{E}_{Q_{O,X}^{0}}\left(\textnormal{IF}(O,X,\psi(Q_{O,X}^{0}))\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\}\right]\right),$
where $\dot{\rho}$ is the derivative of $\rho$ with respect to $\gamma$. Here, $\dot{\rho}$ characterizes the belief on how violations to conditional exchangeability may change the conditional distribution $Q_{O\mid X}^{\gamma}$. For example, when $\rho(O,X,\gamma)=\exp(\gamma\cdot O)$ as in (3), the derivative is $\dot{\rho}(O,X,0)=O$. Therefore, the term $\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\}\right]$ (i.e., $O-\mu(X)$ in (10)) can be viewed as the direction of sensitivity that’s unexplained by the observed variable $X$.
J.3 Extended Remarks from Section 6
@@
Second, we always prioritize the endeavor to meet conditional exchangeability than to make remedies when the conditional exchangeability is violated. Whenever possible, we echo Tipton and Olsen, (2018) and Degtiar and Rose, (2023) in the importance of a careful data collection for the conditional exchangeability to hold, or, for the violation to be as small as possible. In this sense, it’s more preferable to collect more common characteristics in source and target populations, namely a richer $X$. Secondary to that, SLOPE is a useful tool to assess sensitivity/robustness when it is unrealistic to meet the conditional exchangeability with the observed set of $X$. This happens, for example, when (i) it is infeasible to randomize units into the source or target population, (ii) not all covariates can be measured in both populations, which is common if the investigator opts in after data collection, and/or (iii) there exist unobservable differences between populations (e.g., sites) even under a careful design (Allcott,, 2015; Jin et al.,, 2024).
J.4 SLOPE in Other Contexts
@@
J.4.1 SLOPE for Unmeasured Confounding
Let $A\in\{0,1\}$ be the treatment and $Y(A)$ be the potential outcome under treatment $A$, where $A=1$ means intervention and $A=0$ means control. We suppose SUTVA (Assumption 4) holds; i.e., when $A=a$, the observed outcome $Y$ is $Y(a)$, for $a=0,1$. Suppose we are interested in the average treatment effect on the treated, a popular estimand in causal inference,
@@
where we use $\cdot\mid X,A]$ to represent conditional densities (provided exist with respect to some measure) on the underlying population. For more explanations on the sensitivity model (67), see Scharfstein et al., (1999); Franks et al., (2020).
Finally, we make a clarification on this setting. The current section should not be confused with Section D. In Section D, we transport treatment effect from a source population to a target population where we have assumed that the treatment effect can be identified within the source; in other words, the “causal” assumptions already hold in the source population. However, in this section, we no longer consider any transfer learning setting but instead focus on a causal setting with potential unmeasured confounding. Here, may still think of $P$ (i.e., population of control units) as the “source” population and $Q$ (i.e., population of treated units) as the “target” population, bearing in mind the difference in the meaning of conditional exchangeability assumption as well as the different interpretations of the corresponding SLOPEs.
J.4.2 SLOPE for Non-ignorable Missing
@@
J.5 Remark on Extending SLOPE to Bound-Based Sensitivity Analysis Models
@@
We consider a simplified version of the setting in Zeng et al., (2023) and their sensitivity model. Suppose the target estimand is the mean, $\psi^{{\textnormal{mean}}}(Q_{O,X})=\textnormal{E}_{Q_{O,X}}(O)$. Let $\gamma$ be the sensitivity parameter and suppose the target conditional distribution $Q_{O\mid X}^{\textnormal{bias}}$ deviates from the source conditional distribution $P_{O\mid X}$ so that the bias in the conditional mean is no larger than $\gamma$:
Consequently, the target estimand can be bounded as
@@
J.6 Connections to Marginal Interventional Effect
@@
We begin with a brief introduction to the marginal interventional treatment effect; see Zhou and Opacic, (2022) for details. In causal inference (see notation in Section D.1), an interventional effect quantifies the change in outcome when the propensity of a proportion of units changes. This is in contrast to conventional estimands, like the average treatment effect, which compare the outcome when all units receive treatment with the outcome when all receive placebo. To define an interventional effect, suppose the propensity score is $\pi(Z)=P(A=1\mid Z)$ and consider changing the propensity in a way of $\pi^{\gamma}(Z)$ as a function of $\pi(Z)$ with $\pi^{0}(X)=\pi(Z)$. Then under assumptions in Section 3 of Zhou and Opacic, (2022), the interventional effect (IE) is
@@
The MIE with IPSI is connected to SLOPE for two reasons. First, the IPSI shifts the intervention $\pi^{\gamma}(Z)$ in the same exponential tilting model as in the sensitivity model (3); see below for details. Second, the concept of MIE, which defines a treatment effect in terms of the change in the outcome with an infinitesimal change in propensity, is broadly connected to the concept of SLOPE. To see the connections more explicitly, we next re-express the SLOPE under the regime of incremental propensity score intervention. We note that although the resulting quantity no longer represents a “‘sensitivity to local perturbation from the exchangeability”, we still refer to it as SLOPE for ease of communication, keeping in mind that we temporarily treat SLOPE as a mathematical quantity instead of a sensitivity/robustness measure.
Let $O=A$ be the binary treatment. Suppose $P$ is the population of interest and $P_{Z,A,Y}$ is the joint distribution of pre-treatment covariate $Z$, binary treatment $A$ and outcome $Y$. We adopt the potential outcome framework (Section D.1) and suppose SUTVA (Assumption 4) and strong ignorability (Condition 2) hold.

### v2

@@ I.2.3 Proof of Proposition 2
This section details some extended remarks deferred from the main text. Section J.1 discusses vector valued SLOPE. Section J.2 includes some extended remarks from Section 6. Section J.3 discusses defining SLOPE for other types of conditional exchangeability assumptions, including the no unmeasured confounding assumption and the missing at random assumption. Section J.4 details Remark 2 on the challenge of extending the notion of SLOPE to other bound-based sensitivity models. Section J.5 states the mathematical connection between SLOPE and the marginal interventional effect (Zhou and Opacic,, 2022) with the incremental propensity score intervention (Kennedy,, 2019).
@@
When the functional $\psi(\cdot)$ is vector valued with dimension $p$, the corresponding SLOPE, as defined in Definition 1, is also a vector of $p$ elements. Each dimension of SLOPE represents the robustness of the corresponding element of the target estimand when conditional exchangeability is “near” violated. For example, when $X$ is $p$-dimensional and the target estimand is a $p$-dimensional vector of ordinary least squares (OLS) coefficients of regressing $O$ on $X$, then each element of SLOPE describes the robustness of the corresponding coefficient; see Section C.5 for a formal presentation of the SLOPE and Section E.5 for the estimation. More generally, for SLOPE as vectors, the connection between SLOPE and IF still holds (i.e., Theorem 3). In addition, the two proposed estimators are still applicable, as mentioned in Remarks 4 and 5.
The main difference between a vector of SLOPE and a scalar SLOPE is on the interpretation, and in particular, the informed guidance in designs. Specifically, as mentioned in Section 3.1, when SLOPE becomes a vector and all elements are of scientific interest, then practitioners need to find an appropriate summary of the vector in order to compare SLOPEs across study designs. This could be a visual summary, such as plotting the SLOPE of each dimension, or a quantitative measure such as a norm of the SLOPE.
Example 6 (SLOPE for OLS Coefficient).
Suppose $X$ is a $p$-dimensional vector and $O$ is an outcome variable of interest. The target estimand is defined as the regression coefficients of regressing $O$ on $X$. In other words, consider $\psi^{{\textnormal{OLS}}}$ such that
$\textnormal{E}_{{\textnormal{OLS}}}\left(XX^{\intercal}\psi^{{\textnormal{OLS}}}-XO\right)=0,$
which we suppose the uniqueness of $\psi^{{\textnormal{OLS}}}(Q_{O,X})$. Following Section C.5, we know the SLOPE for $\psi^{{\textnormal{OLS}}}$ is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$ $=\left\{\textnormal{E}(XX^{\intercal})\right\}^{-1}\textnormal{E}_{Q_{X}}\left\{\sigma^{2}(X)X\right\},$
which is a $p$-dimensional vector.
To compare SLOPE across designs, we define the magnitude of SLOPE in two scenarios. First, suppose $X$ has been standardized within the source and the target population. Then we can define the norm of SLOPE as the $L_{2}$ norm of $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})$. With this notion of the magnitude, interpretations of SLOPE based on its magnitude that were discussed in Section 4.1 are applicable. Second, suppose $X$ has not been standardized, then we consider the Mahalanobis distance with covariance matrix $\{\textnormal{E}_{Q_{X}}(XX^{\intercal})\}^{-1}$. More specifically, we define the magnitude of SLOPE as
$\left\{\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})\right\}^{\intercal}\textnormal{E}_{Q_{X}}(XX^{\intercal})\left\{\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{OLS}}})\right\}$
$=$ $\left[\textnormal{E}_{Q_{X}}\left\{\sigma^{2}(X)X\right\}\right]^{\intercal}\left\{\textnormal{E}_{Q_{X}}(XX^{\intercal})\right\}^{-1}\textnormal{E}_{Q_{X}}\left\{\sigma^{2}(X)X\right\}.$
In practice, the SLOPE needs to be estimated. Example estimators of the SLOPE for OLS coefficients are provided Section E.5.
J.2 Extended Remarks from Section 6
@@
Second, as mentioned in the main text, we echo Tipton and Olsen, (2018) and Degtiar and Rose, (2023) in the importance of a careful data collection for the conditional exchangeability to hold, or, for the violation to be as small as possible. In this sense, it’s more preferable to collect more common characteristics in source and target populations, namely a richer $X$. Secondary to that, SLOPE is a useful tool to assess sensitivity/robustness when it is unrealistic to meet the conditional exchangeability with the observed set of $X$. This happens, for example, when (i) it is infeasible to randomize units into the source or target population, (ii) not all covariates can be measured in both populations, which is common if the investigator opts in after data collection, and/or (iii) there exist unobservable differences between populations (e.g., sites) even under a careful design (Allcott,, 2015; Jin et al.,, 2024).
J.3 SLOPE in Other Contexts
@@
J.3.1 SLOPE for Unmeasured Confounding
Let $A\in\{0,1\}$ be the treatment and $Y(A)$ be the potential outcome under treatment $A$, where $A=1$ means intervention and $A=0$ means control. We suppose SUTVA (Assumption 4) holds; i.e., when $A=a$, the observed outcome $Y$ is $Y(a)$, for $a=0,1$. To fix ideas, suppose we are interested in the average treatment effect on the treated, a popular estimand in causal inference,
@@
where we use $\cdot\mid X,A]$ to represent conditional densities (provided exist with respect to some measure) on the underlying population. For more explanations on the sensitivity model (68), see Scharfstein et al., (1999) and Franks et al., (2020).
Finally, we make a clarification on this setting. The current section should not be confused with Section D. In Section D, we transport treatment effect from a source population to a target population where we have assumed that the treatment effect can be identified within the source; in other words, the “causal” assumptions already hold in the source population. However, in this section, we no longer consider any transfer learning setting but instead focus on a causal setting with potential unmeasured confounding. Here, one may still think of $P$ (i.e., population of control units) as the “source” population and $Q$ (i.e., population of treated units) as the “target” population, while keeping in mind the difference in the meaning of conditional exchangeability assumption as well as the different interpretations of the corresponding SLOPEs.
J.3.2 SLOPE for Non-Ignorable Missingness
@@
J.4 Remark on Extending SLOPE to Bound-Based Sensitivity Analysis Models
This section provides details of Remark 2 in the main text.
@@
We consider a simplified version of the setting in Zeng et al., (2023) and their sensitivity model. Suppose the target estimand is the mean, $\psi^{{\textnormal{mean}}}(Q_{O,X})=\textnormal{E}_{Q_{O,X}}(O)$. Let $\gamma$ be the sensitivity parameter and suppose the target conditional distribution $Q_{O\mid X}^{\textnormal{bias}}$ deviates from the source conditional distribution $P_{O\mid X}$ so that the bias in the conditional mean is no larger than $\gamma$, as sated in (15). Consequently, the target estimand can be bounded as
@@
J.5 Connections to Marginal Interventional Effect
@@
We begin with a brief introduction to the marginal interventional treatment effect; see Zhou and Opacic, (2022) for details. In causal inference (see notation in Section D.1), an interventional effect quantifies the change in outcome when the propensity of a proportion of units changes. This is in contrast to conventional estimands, like the average treatment effect and the quantile treatment effect, which compare the outcome when all units receive treatment with the outcome when all receive placebo. To define an interventional effect, suppose the propensity score is $\pi(Z)=P(A=1\mid Z)$ and consider changing the propensity in a way of $\pi^{\gamma}(Z)$ as a function of $\pi(Z)$ with $\pi^{0}(X)=\pi(Z)$. Then under assumptions in Section 3 of Zhou and Opacic, (2022), the interventional effect (IE) is
@@
The MIE with IPSI is connected to SLOPE for two reasons. First, the IPSI shifts the intervention $\pi^{\gamma}(Z)$ in the same exponential tilting model as in the sensitivity model (3); see below for details. Second, the concept of MIE, which defines a treatment effect in terms of the change in the outcome with an infinitesimal change in propensity, is broadly connected to the concept of SLOPE. To see the connections more explicitly, we next re-express the SLOPE under the regime of incremental propensity score intervention. We note that although the resulting quantity no longer represents the “sensitivity to local perturbation from the exchangeability”, we still refer to it as SLOPE for ease of communication, keeping in mind that we temporarily treat SLOPE as a mathematical quantity instead of a sensitivity/robustness measure.
Let $O=A$ be the binary treatment. Suppose $P$ is the population of interest and $P_{Z,A,Y}$ is the joint distribution of pre-treatment covariate $Z$, binary treatment $A$ and outcome $Y$. We adopt the potential outcome framework (Section D.1) and suppose SUTVA (Assumption 4) and strong ignorability (Assumption 3) hold.